### v1

@@ 0.A.2 Selecting Annotators
It is possible that, being a somewhat homogeneous group, annotators may share a similar perception of what is considered normal or abnormal in surveillance camera videos.
@@
The purpose of this test is solely to evaluate our pre-annotation pipeline. Although the potential bias introduced into the data by pre-annotations is studied in 4.2.
0.A.3 Platform and Infrastructure Choices
@@
We used Label Studio [24] for our annotation needs.
@@
0.A.4 Dataset Analysis and Evaluation
@@
An extensive study on the consistency of annotations and, consequently, the patterns of disagreement has been conducted; it is detailed in [citation].

### v2

@@ B.2 Selecting Annotators
It is possible that, being a somewhat homogeneous group, annotators may share a similar perception of what is considered normal or abnormalin the videos.
@@
The purpose of this test is solely to evaluate our Pre-Annotation pipeline. Although the potential bias introduced into the data by Pre-Annotations is studied in Section 4.
B.3 Platform and Infrastructure Choices
@@
We used Label Studio (Tkachenko et al., 2020) for our annotation needs.
@@
B.4 Dataset Analysis and Evaluation
@@
An extensive study on the consistency of annotations and, consequently, the patterns of disagreement has been conducted; it is detailed in Section 4.